1 liter of air at 76cm of mercury is compressed to a pressure of 152cm of mercury under isothermal conditions .What is new volume?

Grade:12th pass

2 Answers

karl
6 years ago
Under isothermal condition PV=constant or,P1V1=P2V2 substituting the given values we have 760*1=1520*V2 V2=760/1520 =0.5 L
vamsi
6 years ago
arrey dont you know even this question, this is very simple for those who passed 10th class cause it is also related to gaseous state the answer is p1v1=p2v2 substute you will get answer